The pressure applied by a boy on a bicycle pump was 2 Pascal. If the volume was increased to 2/3 the initial volume. Calculate the final pressure.


1
Expert's answer
2021-05-24T07:22:17-0400

Initial pressure, "P_1=2\space Pa"

Initial volume, "V_1=V"

Final pressure, "P_2=P"

Final Volume, "V_2=V+\dfrac{2}{3}V=\dfrac{5}{3}V"

"P_1V_1=P_2V_2"

"P=\dfrac{P_1V_1}{V_2}=\dfrac{2\times V}{\dfrac{5}{3}V}"

"P=1.2\space Pa"